OpenStudy (anonymous):

A high-altitude spherical weather balloon expands as it rises due to the drop in atmospheric pressure. Suppose that the radius r increases at the rate of 0.07 inches per second and that r = 36 inches at time t = 0. Determine the equation that models the volume V of the balloon at time t and find the volume when t = 400 seconds.

OpenStudy (anonymous):

Calc?

OpenStudy (anonymous):

Yeah

OpenStudy (anonymous):

this is a related rates problem. you have volume which depends on radius but radius is some other function of time (perhaps). so when you take the derivative of volume, you have to apply the chain rule. \[V=\frac{ 4 }{ 3 }\pi r ^{3} \text{, } \frac{ dV }{ dt }=4\pi r ^{2}\frac{ dr }{ dt }\] does this make sense so far?

OpenStudy (anonymous):

Kind of what is the dV/dt but the V=4 /3 pi r ^3 is the formula

OpenStudy (anonymous):

oops... i mis-read. they just want you to put in the radius function into the volume formula. can you tell me what the radius should be as a function of time? let's call that r(t).

OpenStudy (anonymous):

So it would be 36(0.07)

OpenStudy (anonymous):

not exactly... where does t (time) fit into it?

OpenStudy (anonymous):

what would it be when t is 0? when t is 1? etc...

OpenStudy (anonymous):

when t is 0 it would be 0,07 right?

OpenStudy (anonymous):

Suppose that the radius r increases at the rate of 0.07 inches per second and that r = 36 inches at time t = 0. so when t is 0, r = ?

OpenStudy (anonymous):

36 ?

OpenStudy (anonymous):

yep. then what about when t is 1?

OpenStudy (anonymous):

Would it be 1* 0.07 then multiply it by 36 ?

OpenStudy (anonymous):

nope... let's read Suppose that the radius r increases at the rate of 0.07 inches per second and that r = 36 inches at time t = 0. so it's increasing by .07 inches per second. 36 + .07t where t is in seconds. so after 1 second it would be 36+.07*1 = 36+.07 = 36.07 inches make sense?

OpenStudy (anonymous):

Ok yeah now it does.

OpenStudy (anonymous):

so r(t) = 36 + .07t then\[V \left( t \right)=\frac{ 4 }{ 3 }\pi \left[ r \left( t \right) \right]^{3}\] does that make sense?

OpenStudy (anonymous):

Yeah

OpenStudy (anonymous):

just plug in what r(t) is...

OpenStudy (anonymous):

Ok thanks could you help with an other problem

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

you want to post here or as a new problem?

OpenStudy (anonymous):

Ok let me post it as new problem